1. A uniform electric field with a magnitude of 6.8 x 104 N/C is passing through a flat square with a total area of area 600 cm. Calculate the electric flux.


2. A uniform electric field 5000 N/C passing through a flat square area A = 2.75 m2. Determine the electric flux.


3. A 9.8 x 10 -9-point charge is at the center of a sphere 0.75 m in diameter. The area of sphere can be calculated using A= 4πr2. Determine the (a) the magnitude of the electric field at any point on its surface, (b) the angle that the electric field at a point on the surface with respect to the normal to the surface at that point, and (c) the electric flux that passes through the entire surface of the sphere.


1
Expert's answer
2021-04-28T20:18:48-0400

1. The flux is


"\Phi_E=EA=(6.8·10^4)(600·10^{-6})=40.8\text{ Nm}^2/\text C."

2. The flus is


"\Phi_E=EA=(5000)(2.75)=13750\text{ Nm}^2/\text C."

3. (a) The field on the surface of the sphere:


"E=\frac{kq}{r^2}=626\text{ N/C}."

(b) The angle that the electric field vector at a point on the surface with respect to the normal to the surface at that point makes is 0°.

(c) The electric flux that passes through the entire surface of the sphere is


"\Phi_E=EA=E\pi r^2=\\=(626)(3.14·0.375^2)=277\text{ Nm}^2/\text C."